OpenStudy (anonymous):

lim of (x-2)/(x^2 -x-2) as x approaches 2?? idk how to solve limits...

OpenStudy (anonymous):

first check that you get \(\frac{0}{0}\) when you replace \(x\) by \(2\) you do so 1) factor 2) cancel 2) replace \(x\) by 2

OpenStudy (anonymous):

\[\frac{x-2}{x^2-x-2}=\frac{x-2}{(x-2)(x+1)}=\frac{1}{x+1}\] is the calculation don't think too hard about how to factor, because you know if \(2\) gets you 0, then \(x-2\) must be a factor now let \(x=2\)

OpenStudy (anonymous):

then it's zero?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh for the last one? 1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does thatmean that the limit is 1/3?

OpenStudy (anonymous):

wait so is a limit always a y value

OpenStudy (anonymous):

yes and yes